OpenStudy (lgbasallote):

help? \(\LARGE \int \frac{\ln x -4}{x(\ln ^2 x + 3)} dx\) show steps pleeeeaaasssee...i think it has something to do with the inverse trig functions

OpenStudy (lgbasallote):

@Callisto ? you're the only one online i know that knows calculus lol

OpenStudy (lgbasallote):

@Mimi_x3 !!! help :P

OpenStudy (mimi_x3):

Try let u=logx

OpenStudy (mimi_x3):

u=lnx in this case.

OpenStudy (lgbasallote):

\(\LARGE \int \frac{u-4}{u^2 + 3} du\)

OpenStudy (mimi_x3):

break it up into two fractions.

OpenStudy (lgbasallote):

\(\LARGE \int \frac{u}{u^2+3}du - \int \frac{4}{u^2 + 3}du\)

OpenStudy (mimi_x3):

Yes! The first one is a substitution; and the next is a trig inverse.

OpenStudy (mimi_x3):

Are you able to do it?

OpenStudy (lgbasallote):

let me see...

OpenStudy (mimi_x3):

well, they are both easy; if need help just ask me. xD

OpenStudy (lgbasallote):

easy for you :P i still need practice to see these things easier

OpenStudy (mimi_x3):

okay sorry..well; just try it, you can do it!

OpenStudy (lgbasallote):

\(\LARGE \ln (\sqrt{u^2 +3}) - \frac{4\sqrt 3 \tan^{-1} u}{3\sqrt 3}\)

OpenStudy (lgbasallote):

lol ignore the \(\sqrt 3\)'s forgot to cancel them

OpenStudy (lgbasallote):

am i right @Mimi_x3 ?

OpenStudy (mimi_x3):

wait..something looks wrong..

OpenStudy (lgbasallote):

:O

OpenStudy (mimi_x3):

Congratulations! Correct! ^_^

OpenStudy (mimi_x3):

don't forget to sub back the x!

OpenStudy (lgbasallote):

yayyy!!! but how come when i input the answer in wolfram and took the derivative the second term says 4/3(u^2 +1) is that the same?

OpenStudy (mimi_x3):

why did you take the derivative...?

OpenStudy (lgbasallote):

to see if it will come out as the integral

OpenStudy (lgbasallote):

because taking the integral of that from the start results in crazy manipulations

OpenStudy (mimi_x3):

i don't think that it will work since it's a trig inverse..

OpenStudy (mimi_x3):

wait..which integral are you talking about \[\int\limits\frac{4}{x^{2}+3} dx\]??

OpenStudy (lgbasallote):

oh wait lol it should be \(\LARGE \frac{4 \tan^-1 (\frac {u}{sqrt 3})}{sqrt 3}\)

OpenStudy (lgbasallote):

wonder where i went wrong

OpenStudy (mimi_x3):

you did not do it wrong..

OpenStudy (mimi_x3):

your answer was like that..

OpenStudy (lgbasallote):

no...my sqrt 3 was not in the argument...

OpenStudy (lgbasallote):

hah lol found it...forgot to put over sqrt 3 in the argument lol

OpenStudy (lgbasallote):

it was a simple typo =))

OpenStudy (mimi_x3):

lol, i was going to type the solution up.. xD So need anymore help?

OpenStudy (lgbasallote):

yeah :DD just some tips though....

OpenStudy (mimi_x3):

Okay; what tips..?

OpenStudy (lgbasallote):

i meant hints while solving...ill post another problem

OpenStudy (mimi_x3):

alrighty! try me; too tired now..